The sum of three consecutive even integers is less than or equal to 42. Find the largest se of these numbers.
Answer by stanbon(75887)

The sum of three consecutive even integers is less than or equal to 42. Find the largest se of these numbers.
EQUATION:
x + (x+1) + (x+2) = 42
3x+3 = 42
3x = 39
x = 13
x+1=14
x+2=15
================
Cheers,
Stan H.